# UniREditBench: A Unified Reasoning-based Image Editing Benchmark

### v1

@@ 5.1 Implementation Details
Baselines. We benchmark closed-source models including: GPT-4o [12], Nano Banana [7], Gemini-2.0 [5], Seedream4.0 [24], Wan 2.5 [28], and FLUX-Kontext-Pro [2], as well as open-source models including: Bagel [6], Qwen-Image-Edit [39], Step1X-Edit [20], FLUX.1-Kontext-dev [2], Emu2 [26], Omnigen2 [40], Omnigen [42], HiDream-Edit [4], MagicBrush [50], and AnyEdit [47].
Training and Evaluation. We train all components of Bagel [6] except the VAE model for 5,000 iterations using the Adam optimizer and the cosine learning rate scheduler on UniREdit-Data-100K. The scheduler includes 500 warm-up steps, with the peak learning rate of 2e-5 and minimum learning rate of 1e-6. During inference, we use the official inference settings provided by Bagel. To ensure fair comparisons with other baselines, we adopt the original inference configurations of these models.
@@ 5.2 Benchmarking Results on UniREditBench
As shown in Tab. 3.3, among closed-source models, GPT-4o achieves the highest average performance across all scenarios, with Nano Banana showing comparable capability. Wan2.5 delivers balanced results on real-world tasks but lags on game scenarios that require strategic reasoning. Besides, Seedream4.0 performs reasonably well on the structure transform dimension yet encounters challenges in game scenarios. Among open-source baselines, Qwen-Image-Edit performs strongly on real-world tasks such as attribute modification and structure transform. However, most models remain comparatively weak on game scenarios like Strategic Reasoning. Overall, compared with open-source methods, closed-source models, particularly GPT-4o, maintain a clear advantage. While some open-source models are competitive on specific real-world tasks, they generally struggle with complex reasoning in game scenarios. Notably, only GPT-4o and Nano Banana achieve an average score greater than 60 on game scenarios, underscoring that this setting remains highly challenging and serves as a useful test for current models.
@@ 5.3 Comparison Results of UniREdit-Bagel
Table 3: Out-of-distribution quantitative performance comparison on RISEBench [52]. GPT-4.1 is used as the evaluator. Best scores are in bold.
@@ 6 Conclusion
This paper introduces UniREditBench, a unified reasoning-based benchmark for image editing with broader evaluation dimension coverage and a reliable dual-reference evaluation pipeline. Additionally, we design a multi-scenario data synthesis pipeline and release UniREdit-Data-100K, a large-scale dataset with high-quality chain-of-thought (CoT) annotations. To demonstrate its effectiveness, we fine-tune Bagel on this dataset, resulting in UniREdit-Bagel, which achieves significant improvements both quantitatively and qualitatively. Through comprehensive benchmarking of both open-source and closed-source image editing models, we highlight their strengths and weaknesses across various aspects.
@@ A.1 Instruction De-duplication
  • Semantic-Similarity Deduplication: We use a sentence-transformers [23] model to extract sentence embeddings for both the original image description and edit instruction. We then compute the pairwise similarity between these embeddings. If the similarity score exceeds a threshold of 0.7 for either description and instruction, the samples are deemed semantically redundant and are filtered out to enhance dataset diversity.
@@ A.2 Quality Filtering
Table 4: Quantitative comparisons on KRISBench [41]. GPT-4.1 is used as the evaluator. Best scores are in bold while second-best is underlined.
@@
FLUX-Kontext-Pro 34.75 56.62 54.95 36.36 42.15 42.52 43.89 38.01 42.33 43.54 47.45 34.09 50.55 34.88 58.44 26.88 46.70 52.90 43.72
Seedream4.0 64.77 74.32 80.19 66.07 59.48 64.68 79.48 61.58 63.21 39.27 46.28 43.00 41.34 48.55 38.43 68.75 54.15 33.43 57.05
Wan2.5 72.97 75.30 79.97 59.86 64.51 66.56 61.60 63.67 66.66 63.39 44.67 53.73 40.00 58.54 60.59 45.94 65.17 47.47 60.59
Nano Banana 75.37 79.72 85.55 71.35 70.65 73.36 73.22 79.44 75.34 66.26 61.43 54.65 54.40 64.65 40.92 91.99 62.65 50.05 68.38
GPT-4o 83.55 82.98 92.16 75.70 76.10 76.32 88.88 78.97 76.28 78.82 58.88 44.95 47.02 64.55 63.80 68.62 82.27 49.77 71.64
@@
MagicBrush 35.65 50.55 48.34 44.13 45.07 47.83 38.73 47.88 47.52 63.43 44.74 26.26 30.30 34.12 31.05 40.33 32.65 29.00 40.98
Omnigen2 50.32 58.03 67.73 47.77 49.90 57.63 51.09 55.05 51.98 70.78 51.17 1.50 30.17 23.07 46.17 4.85 39.63 34.58 43.96
Step-1X-Edit 51.82 67.18 63.29 51.45 54.98 62.18 51.03 55.72 57.78 62.90 33.62 35.40 33.48 43.02 49.92 39.17 54.53 34.73 50.12
Bagel-Think 58.38 64.15 63.38 56.27 54.40 58.63 48.82 56.47 57.75 65.65 47.30 42.12 40.97 47.80 40.35 32.40 48.83 38.80 51.25
Qwen-Image-Edit 72.08 78.31 81.75 64.57 68.78 69.51 77.12 67.83 67.25 57.03 36.58 37.79 30.02 48.47 33.00 31.63 60.02 34.71 56.46
UniREdit-Bagel (Ours) 84.43 74.06 85.13 72.08 74.31 71.95 83.05 71.17 72.45 84.93 86.93 61.25 73.76 87.48 70.65 93.90 97.73 74.55 78.87
@@ A.3 Human Inspection
Two web interfaces are shown in Figs. 11 and  12.
@@ Appendix D More Qualitative Comparison Results
We provide additional qualitative comparisons on UniREditBench in Fig. 8 and 9, and comparisons on RISEBench in Fig. 10.
@@
Appendix E Ethical statement

### v2

@@ 5.1 Implementation Details
Baselines. We benchmark closed-source models including: GPT-4o [hurst2024gpt], Nano Banana [Gemini-2.5-Flash-Image], Gemini-2.0 [comanici2025gemini], Seedream4.0 [seedream4], Wan 2.5 [wan2025wan], and FLUX-Kontext-Pro [flux_kontext], as well as open-source models including: Bagel [deng2025bagel], Qwen-Image-Edit [qwen-image], Step1X-Edit [liu2025step1x], FLUX.1-Kontext-dev [flux_kontext], Emu2 [sun2024generative], Omnigen2 [wu2025omnigen2], Omnigen [xiao2025omnigen], HiDream-Edit [cai2025hidream], MagicBrush [zhang2023magicbrush], and AnyEdit [yu2025anyedit].
Training and Evaluation. We train all Bagel [deng2025bagel] components, except the VAE, for 5,000 iterations on UniREdit-Data-100K using the Adam optimizer and a cosine learning-rate schedule with 500 warm-up steps, a peak learning rate of $2\times 10^{-5}$, and a minimum learning rate of $10^{-6}$. The loss weights are set to $\lambda_{\text{text}}=2$ and $\lambda_{\text{img}}=1$. During inference, we use the official inference settings provided by Bagel. To ensure fair comparisons with other baselines, we adopt the original inference configurations of these models.
@@ 5.2 Benchmarking Results on UniREditBench
As shown in Tab. 3.3, among closed-source models, GPT-4o achieves the highest average performance across all scenarios, with Nano Banana performing comparably. Wan2.5 delivers balanced results on real-world tasks but lags on game scenarios that require strategic reasoning. Besides, Seedream4.0 is competitive on structure transform yet encounters challenges in game scenarios. Among open-source baselines, Qwen-Image-Edit performs strongly on real-world tasks such as attribute modification and structure transform. However, most models remain comparatively weak on game scenarios like Strategic Reasoning. Overall, compared with open-source methods, closed-source models, particularly GPT-4o, maintain a clear advantage. While some open-source models are competitive on specific real-world tasks, they generally struggle with complex reasoning in game scenarios. Notably, only GPT-4o and Nano Banana achieve an average score greater than 60 on game scenarios, underscoring that this setting remains highly challenging and serves as a useful test for current models.
@@ 5.3 Comparison Results of UniREdit-Bagel
Table 3: Out-of-distribution quantitative performance comparison on RISEBench [zhao2025envisioning]. GPT-4.1 is used as the evaluator. Best scores are in bold.
@@ 6 Conclusion
This paper presents UniREditBench, a unified reasoning-based benchmark for image editing with broad dimension coverage and a robust dual-reference evaluation protocol. We further introduce a multi-scenario synthesis pipeline and release UniREdit-Data-100K, a large-scale dataset with high-quality CoT annotations. To demonstrate its effectiveness, we fine-tune Bagel on this dataset and yields UniREdit-Bagel, which achieves substantial quantitative and qualitative gains. Comprehensive benchmarking of open-source and closed-source image editing models reveals their strengths and weaknesses across various aspects.
@@ A.1 Instruction De-duplication
  • Semantic-Similarity Deduplication: We use a sentence-transformers model to extract sentence embeddings for both the original image description and edit instruction. We then compute the pairwise similarity between these embeddings. If the similarity score exceeds a threshold of 0.7 for either description and instruction, the samples are deemed semantically redundant and are filtered out to enhance dataset diversity.
@@ A.2 Quality Filtering
Table 4: Quantitative comparisons on KRISBench. GPT-4.1 is used as the evaluator. Best scores are in bold while second-best is underlined.
@@
FLUX-Kontext-Pro 37.55 57.15 56.79 37.52 41.46 44.62 47.03 40.56 42.32 49.12 53.18 38.12 54.43 32.45 61.23 27.80 48.45 53.87 45.77
Seedream4.0 64.18 71.78 79.98 65.87 56.10 61.60 77.52 59.15 59.83 38.87 45.47 40.87 40.13 48.05 38.17 68.72 55.27 32.90 55.77
Wan2.5 73.78 75.54 79.95 59.53 62.67 64.38 63.49 60.83 64.91 63.73 43.44 56.57 41.38 58.38 58.57 47.86 64.03 45.97 61.36
Nano Banana 75.82 78.39 86.12 71.63 71.37 71.07 73.14 77.58 71.82 66.74 59.97 54.33 54.04 64.13 39.80 90.81 62.50 51.15 68.26
GPT-4o 83.83 81.52 92.18 77.33 75.86 74.98 88.60 78.98 75.81 77.73 58.96 45.60 49.75 64.62 58.97 67.78 83.73 51.49 73.39
@@
MagicBrush 36.74 51.20 47.75 44.43 44.05 47.06 37.67 45.87 47.42 63.58 45.33 25.60 29.83 33.92 31.45 40.55 33.20 28.23 40.77
Omnigen2 51.25 59.68 68.25 48.28 48.37 54.48 51.48 50.77 50.64 70.28 51.33 1.38 29.80 22.87 44.97 5.10 39.57 32.93 43.41
Lumina-DiMOO 54.22 51.47 54.40 50.23 48.67 52.62 49.65 51.63 50.03 61.23 48.23 36.29 26.35 41.73 65.07 52.47 43.57 35.57 48.54
Step-1X-Edit 52.57 66.82 62.15 50.58 53.07 58.68 49.80 54.23 55.45 65.68 35.05 36.38 33.25 45.49 46.97 39.20 55.94 38.08 50.15
Bagel-Think 59.05 63.84 63.52 55.49 54.05 55.62 48.27 55.03 56.33 66.29 45.48 42.89 41.33 48.42 42.20 33.27 48.62 37.38 50.96
DreamOmni2 61.22 65.82 67.07 49.63 49.58 54.83 53.62 54.11 53.92 72.42 38.78 42.33 45.40 45.86 46.98 51.42 56.55 41.05 52.81
UniWorld-V2 71.45 74.47 83.19 55.55 58.53 65.47 66.95 62.20 61.17 49.27 34.44 42.90 42.65 49.67 34.73 28.20 71.59 35.22 54.87
Qwen-Image-Edit 74.60 76.76 81.45 64.62 66.48 68.78 76.50 64.65 64.69 56.73 38.42 41.02 30.45 48.52 33.60 30.93 61.38 36.21 56.52
UniREdit-Bagel (Ours) 81.08 72.39 84.78 70.81 73.85 72.73 83.13 71.27 71.62 84.90 87.77 57.68 73.05 86.80 71.16 93.20 98.30 71.47 78.15
@@ A.3 Human Inspection
Two web interfaces are shown in Figs. 12 and  13.
@@ Appendix B Detailed Benchmarking Results
Figure 8: Overall performance on UniREditBench of models trained on different size of UniREdit-Data.
@@ Appendix D More Qualitative Comparison Results
We provide additional qualitative comparisons on UniREditBench in Fig. 9 and 10, and comparisons on RISEBench in Fig. 11.
Appendix E Effect of Training Set Size on Overall Score
We provide overall performance comparison on UniREditBench of models trained on difference size of UniREdit-Data, which is visualized in Fig. 8.
@@ Appendix E Effect of Training Set Size on Overall Score
Appendix F Ethical statement